40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 15 10:52:38 2019, gautam, Update, SUS, EY door opened 
    Reply  Tue Jan 15 15:49:47 2019, gautam, Update, SUS, EY door opened EY_OSEMs.png
       Reply  Sat Jan 19 15:33:18 2019, gautam, Update, SUS, ETMY OSEMs faulty Screen_Shot_2019-01-19_at_3.32.35_PM.pngETMY_sensors_1231832635.pdf
          Reply  Tue Jan 22 20:36:53 2019, gautam, Update, SUS, ETMY OSEMs faulty 
             Reply  Wed Jan 23 12:39:18 2019, gautam, Update, SUS, EY chamber work 6x
                Reply  Wed Jan 23 23:12:44 2019, gautam, Update, SUS, Prep for FC cleaning IY_level_before.pdf
                   Reply  Thu Jan 24 15:32:31 2019, gautam, Update, SUS, Y arm cavity side first contact applied Yarm_FC.pdfOSEMfilter.png
                      Reply  Tue Jan 29 22:12:40 2019, gautam, Update, SUS, Alignment prep 
                         Reply  Wed Jan 30 11:54:24 2019, gautam, Update, SUS, More alignment prep IMG_5958.JPGIMG_5962.JPG
                            Reply  Wed Jan 30 19:25:40 2019, gautam, Update, SUS, Xarm cavity alignment Xarm.pngXarm_IR.png
                               Reply  Fri Feb 1 01:24:06 2019, gautam, Update, SUS, Almost ready for pumpdown tomorrow 
                                  Reply  Fri Feb 1 13:16:50 2019, gautam, Update, SUS, Pumpdown 83 underway 
                                     Reply  Fri Feb 1 21:52:57 2019, gautam, Update, SUS, Pumpdown 83 underway 
                                  Reply  Fri Feb 1 14:44:14 2019, gautam, Update, SUS, Y arm FC cleaning and reinstall IMG_5974.JPGIMG_5986.JPGIMG_5992.JPG
    Reply  Wed Jan 16 16:25:25 2019, gautam, Update, SUS, Yarm locked Yarm_locked.png
       Reply  Fri Jan 18 21:34:18 2019, gautam, Update, SUS, Unused optic on EY table IMG_5408.JPG
          Reply  Sat Jan 19 05:07:45 2019, Koji, Update, SUS, Unused optic on EY table 
Message ID: 14422     Entry time: Tue Jan 29 22:12:40 2019     In reply to: 14416     Reply to this: 14423
Author: gautam 
Type: Update 
Category: SUS 
Subject: Alignment prep 

Since we may want to close up tomorrow, I did the following prep work:

  1. Cleaned up Y-end suspension eleoctronics setup, connected the Sat Box back to the flange
    • The OSEMs are just sitting on the table right now, so they are just seeing the fully open voltage
    • Post filter insertion, the four face OSEMs report ~3-4% lower open-voltage values compared to before, which is compatible with the transmission spec for the filters (T>95%)
    • The side OSEM is reporting ~10% lower - perhaps I just didn't put the filter on right, something to be looked at inside the chamber
  2. Suspension watchdog restoration
    • I'd shutdown all the watchdogs during the Satellite box debacle
    • However, I left ITMY, ETMY and SRM tripped as these optics are EQ-stopped / don't have the OSEMs inserted.
  3. Checked IMC alignment
    • After some hand-alignment of the IMC, it was locked, transmission is ~1200 counts which is what I remember it being
  4. Checked X-arm alignment
    • Strictly speaking, this has to be done after setting the Y-arm alignment as that dictates the input pointing of the IMC transmission to the IFO, but I decided to have a quick look nevertheless
    • Surprisingly, ITMX damping isn't working very well it seems - the optic is clearly swinging around a lot, and the shadow sensor RMS voltage is ~10s of mV, whereas for all the other optics, it is ~1mV.
    • I'll try the usual cable squishing voodoo

Rather than try and rush and close up tomorrow, I propose spending the day tomorrow cleaning the peripheral areas of the optic, suspension cage, and chamber. Then on Thursday morning, we can replace the Y-arm optics, try and recover the cavity alignment, and then aim for a Thursday afternoon pumpdown. The main motivation is to reduce the time the optics spend in air after F.C. peeling and going to vacuum.

ELOG V3.1.3-